OpenStudy (anonymous):

(6/5) to the 4 power write your answer as an improper fraction, then as a mixed n number

OpenStudy (ajprincess):

\[(\frac{6}{5})^4\] \[=\frac{6^4}{5^4}\] Can u tell me vat is \(6^4=\)?

OpenStudy (anonymous):

can you type it as a mixed number

OpenStudy (ajprincess):

first can u plz tell me vat is \(6^4\)=? before I come to that part of it.

OpenStudy (anonymous):

24

OpenStudy (ajprincess):

hmm no. \(6^4=6*6*6*6=1296\)

OpenStudy (ajprincess):

Nw can u tell me what is \(5^4=\)?